The sum of ages of john and mark is 29 years.john is 3 years older than mark.how old is john?
Answer by macston(5194)

.
J=John's age; M=Mark's age=J-3
.
J+M=29
J+(J-3)=29
2J=32
J=16
.
ANSWER: John is 16 years old.
.
CHECK:
M=J-3=16-3=13
Mark is 13 years old.
.
J+M=29
16+13=29
29=29
.